OpenStudy (anonymous):

Use the squared identities to simplify 2sin^2sin^2x

OpenStudy (anonymous):

@directrix

OpenStudy (danjs):

u mean sin^2(a) + cos^2(a) = 1 ?

Directrix (directrix):

>> squared identities also known as Pythagorean Identities

OpenStudy (anonymous):

@DanJS I don't know honestly that's how the problem is written and my answer choices are A.3+4cos(2x)+cos(4x)/4 B.3-4cos(2x)+cos(4x)/4 C.3-4cos(2x)-cos(4x)/4 D.3+4cos(2x)-cos(4x)/4

OpenStudy (danjs):

\[2\sin^{2}(\sin^{2}(x)) \]

OpenStudy (danjs):

that what you say?

Directrix (directrix):

I'm attached the "Squared Indentities" here for you to review. And, we'll probably need them here. But, the way you posted this problem, it is hard to tell what is an angle and what we are finding. Use grouping symbols often for clarity.

OpenStudy (anonymous):

The problem is \[2\sin^2x \sin^2x\]

Directrix (directrix):

Okay, now we can get to work.

OpenStudy (danjs):

oh, that makes more sense

OpenStudy (anonymous):

Sorry about that guys

Directrix (directrix):

No worries.

OpenStudy (danjs):

so it is just \[2 \sin^{4}(x) \] not a compound sine

OpenStudy (danjs):

looking at the answers, it seems you need to apply the half angle formulas for sin^2(x)

OpenStudy (danjs):

\[\sin ^{2}x = \frac{ 1-\cos(2x) }{ 2 }\]

OpenStudy (anonymous):

how do I apply the half angle formula? @DanJS

OpenStudy (anonymous):

Do you have any idea how to do this? @Directrix I'm stuck :/

Directrix (directrix):

Did you try taking 2* (sinx)^2 * (sinx)^2 and substituting( (1 - cos(2x) )/2 for each of the (sinx)^2 terms and then multiplying those?

OpenStudy (anonymous):

@Directrix Did you get my message?

Directrix (directrix):

That solution looks right and yields the correct answer. Impressive formatting.

Directrix (directrix):

@Monster11223 ^^

OpenStudy (anonymous):

@Directrix Thanks and it took me a while to do it lol well thats it for now thanks a lot :D

Directrix (directrix):

I'm pleased with your work. See ya later on the OS. Wishing you the best on your exams. Trig always take a lot of time!